OpenStudy (anonymous):

An organic compound consist of A&B elements, if A=68% having atomic mass=22.67&B=32% having atmic mass=5.33 then find its empirical formula?

OpenStudy (anonymous):

(a) AB (b) A2B (c)AB2 (d) N.O.T

OpenStudy (chmvijay):

i did it by trail and error method a) AB total atomic mass =22.67+5.33 = 28 28 gram of total weight contains 5.33 of B then 100 gram of total weight contains = 5.33*100/28 =80.96 not write answer b) A2B =2*22.67+5.33=50.67 50.67 ---5.33 of B 100----5.33 *100 /50.67 = 89.6 c) AB2 =22.67 +2*5.33 =33.33 33.33 --10.66 of B 100 ---10.3*100/ 10.66 ==32 % of B mean 68 % of A hence answer will be C